OpenStudy (anonymous):

Use synthetic substitution to evaluate the polynomial for the given value x = -8. P(x) = x2 - x - 30 P(-8) =

OpenStudy (anonymous):

42

OpenStudy (anonymous):

how, can you explain it to me. Im like struggling with these :/

OpenStudy (anonymous):

replace x with -8: \[(-8)^{2}-(-8)-30=64+8-30=72-30=42\]

OpenStudy (anonymous):

thank you so much! (: